OpenStudy (anonymous):

What is the radius of the circle shown below? x^2+y^2-12x-6y-4=0

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

this is the general equation for circle, the general eqn is this: x^2+y^2+2gx+2fy+c=0 the centre is: (-g, -f), in this case the centre is (-6, 3) the radius is: (g^2+f^2-c)^1/2, in this case the radius is 7.

OpenStudy (anonymous):

(x-h)^2=(y-k)^2=r^2 r=7 k=3 h=6

OpenStudy (anonymous):

the center is 6,3